Laboratory report on determination of unknown resistance by potentialmeter

1
Expert's answer
2022-02-17T11:44:37-0500



Voltage drop across S resistance

"V_s=IR"

Voltage drop across R resistor

"V_R=IS"

Now We know that

"\frac{V_R}{V_S}=\frac{R}{S}"

"R=\frac{V_R}{V_S}\times S"